# Quantformer: from attention to profit with a quantitative transformer trading strategy

### v2

@@ Abstract
In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge. Fully capturing various market variables, including long-term information, as well as essential signals that may lead to profit remains a difficult task for learning algorithms. In order to tackle this challenge, this paper introduces quantformer, an enhanced neural network architecture based on transformers, to build investment factors. By transfer learning from sentiment analysis, quantformer not only exploits its original inherent advantages in capturing long-range dependencies and modeling complex data relationships, but is also able to solve tasks with numerical inputs and accurately forecast future returns over a given period. This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019. The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies. Notably, the model’s innovative use of transformer-liked model to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies. The implementation details and code is available on https://github.com/QuantFormer.
@@
quantformer, transformer, quantitative finance, stock selection, market sentiment
@@ 1 Introduction
The goal of stock trading is to optimize the return on investment in the capital market according to the process of buying or selling one or more companies’ shares. Traders obtain profit when a positive difference is generated by the fluctuation of stock price. However, stocks are influenced by a large number of factors, which constitute a complex system and make it difficult for people to make a profit. The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1]. As a result, many strategies and tools have been built, in parallel with the development of capital markets, and quantitative strategies have been playing an important role among them.
@@
Some traditional quantitative tools, such as the Markowitz portfolio theory [2] and the Capital Asset Pricing Model (CAPM) [3], focus mainly on static fundamental analysis. In other words, these strategies aim to make a profit by simple calculation and analysis. Since then, along with the development of computer science, more quantitative methods and tools have been introduced. Within these methods, factor-based strategies have attracted much attention. In 1993, [4] introduced their Fama-French Three Factor Model (FF3), which has become an influential model in quantitative trading. In 2015, Fama and French revised their model with a Five-Factor Asset Pricing Model (FF5) [5]. Besides this classical theory, numerous trading strategies have been published for decades. [6] built a optimal portfolio to catch future investment opportunities (FIO) by multi-factor models.
@@
In recent years, Machine Learning (ML) has become an instrumental tool for enhancing trading algorithms and decision-making processes. Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets. [7] exploits machine learning algorithms based on manual indicators, but the underlying random walk hypothesis may hamper the task of understanding inherently non-stationary series. With the rise of different architectures, [8] finds some novel approaches to catch the sentiment of the market in trading data, which opened a new field in quantitative trading.
Although there exist several previous experiments with ML factors that attempted to fetch market sentiment in the quantitative finance field, this research still faces two difficulties. Firstly, in the field of sentiment analysis, which is a branch of Natural Language Processing (NLP, a field of computer science that aims to understand, interpret, and generate human language), models are used to convert words in text to word vectors through word embeddings to serve as inputs. However, financial datasets contain categorical data instead of words such as industry types, as well as quantitative data such as price fluctuation, turnover rate, and financial indicators. If the input comprises only categorical data, the time series can be treated as a sentence [9]. In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
Secondly, most NLP tasks can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition. As an example, the transformer architecture is based on the seq2seq architecture [10]. To utilize existing outputs, decoders in transformer sequentially output samples and use masking operations to handle input sequences during training. However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
To address these problems, we propose quantformer, which is a modified transformer architecture adapted to quantitative data, and used as an investment factor. Quantformer is able to input numerical data directly, which refers to a method similar to sentiment analysis. The paper is structured as follows. Section 2 discussed the previous quantitative financial works mainly based on machine learning. Sections 3 and 4 introduce quantformers. A factor based on quantformer will be trained and back-tested. For the practical backtest, we collected data from more than 4,600 stocks in the past 14 years (from 2010 to 2023) from the data-collect platform. To comprehensively test the ability of the factor, we divided the data by different frequencies (Section 4.2) and trained under different training scales (Section 5.3). Finally, the result of the back-tests including the comparison between the quantformer-factor and other 100 factors as well as the insights gained from such comparative analysis will be discussed in Sections 4 and 5.
@@ 2 Preliminaries
This section briefly introduces the related works about stock prediction with market sentiment and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
@@ 2.1 Stock data and market sentiment
It is known that, in some cases, stock data can be deemed to influence the future trend of a stock price [11]. [12, 13] have discussed the correlation between investor sentiment and stock market metrics such as turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns. This relationship is not uniform but varies with changes in investor sentiment. It also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions. This link between sentiment and market performance is further affirmed by studies focusing on factors such as the turnover ratio, which are seen as reflective of market liquidity and investor behavior [14]. [15] built a new factor based on sentiment analysis with the average of trading signals from technical trading strategies to benchmark stocks of the S&P 500 index and DJIA. The sentiment factor shows correlation with stock returns. These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
@@ 2.2 Applications of ML methods for trading
2.2.1 Support Vector Machines (SVM)
SVM has been used in quantitative trading, as it can manage high-dimensional spaces and intricacies found in datasets. [16] used SVMs to forecast financial time series in 2015, with a focus on the South Korean stock market, exhibiting the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting. Similarly, [17] showcased SVMs’ prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments. [18] assessed SVMs’ adaptability and precision in financial tasks such as portfolio and quantitative trading.
However, SVMs can also show some disadvantages in quantitative trading. In [19], SVM was underscored for the accuracy of overfitting. When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, making the strategy ineffective on unseen data. In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptations [20].
@@ 2.2.2 Long Short-Term Memory (LSTM)
LSTM is a neural network architecture designed to capture long-term dependencies in sequential data [21]. LSTM neural networks can accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models. [22] combined LSTM with linear Kalman filter and set experiments on different volatility stocks to build strategies. [23] deployed LSTM with stacked autoencoders to forecast financial time series. [24] tested the efficacy of LSTM for price predictions they observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over other conventional models. [25] combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy. For predicting stock prices, [26] harnessed the power of LSTM to unearth multi-frequency trading patterns. Their method underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
@@ 2.2.3 Gated Recurrent Units (GRUs)
GRUs [27] are another neural network architecture which has emerged as a method in the quest for accurate financial forecasting. GRUs can capture long-term dependencies in time series data and have seen increasing popularity in the world of quantitative trading. [23] introduced a comprehensive deep-learning framework for financial time series, which incorporated GRUs alongside LSTM and stacked autoencoders. Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs. [28] centralized on LSTM networks, provided comparative insights into GRUs. However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints. Beyond that, [29] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [30]. The strategy to combine CNNs and GRUs also exhibits advantages for stock price prediction. [31] introduced the Attention-GRU model which combined attention and GRU to establish a new factor based on CVaR [32, 33] portfolio. The Attention-GRU model fitted market return, stocks return from 28 Dow Jones Industrial Average index (DJIA) stocks, which achieved better performance than other models in 8 metrics such as annual return, standard deviation and information ratio.

### v3

@@ Abstract
In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge. Fully capturing various market variables, including long-term information, as well as essential signals that may lead to profit remains a difficult task for learning algorithms. In order to tackle this challenge, this paper introduces quantformer, an enhanced neural network architecture based on transformer, to build investment factors. By transfer learning from sentiment analysis, quantformer not only exploits its original inherent advantages in capturing long-range dependencies and modeling complex data relationships, but is also able to solve tasks with numerical inputs and accurately forecast future returns over a given period. This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2023. The results of this study demonstrate the model’s superior performance in predicting stock trends compared with other 100-factor-based quantitative strategies. Notably, the model’s innovative use of transformer-like model to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies. The implementation details and code is available on Github.
@@
quantformer, transformer, neural networks, quantitative finance, stock selection, portfolio optimization, market sentiment
journal: arXiv
[…]
@@ 1 Introduction
The goal of stock trading is to optimize the return on investment in the capital market according to the process of buying or selling one or more companies’ shares. Traders obtain profit when a positive difference is generated by the fluctuation of stock price. However, stocks are influenced by a large number of factors, which constitute a complex system and make it difficult for people to make a profit. The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling (Wang et al., 2022b). As a result, many strategies and tools have been built, in parallel with the development of capital markets, and quantitative strategies have been playing an important role among them.
Some traditional quantitative tools, such as the Markowitz portfolio theory (Markowitz, 1952) and the Capital Asset Pricing Model (CAPM) (Sharpe, 1964), focus mainly on static fundamental analysis. In other words, these strategies aim to make a profit by simple calculation and analysis. Since then, along with the development of computer science, more quantitative methods and tools have been introduced. Within these methods, factor-based strategies have attracted much attention. In 1993, Fama and French (1993) introduced the Fama-French Three Factor Model (FF3), which has become an influential model in quantitative trading. In 2015, Fama and French revised their model with a Five-Factor Asset Pricing Model (FF5) (Fama and French, 2015). Besides this classical theory, numerous trading strategies have been published for decades. For example, Shi et al. (2024) built an optimal portfolio to catch Future Investment Opportunities (FIO) by multi-factor models.
@@
In recent years, Machine Learning (ML) has became an instrumental tool in trading algorithms and decision-making processes. ML methods allow systems to learn from and make decisions based on data, and lend themselves particularly well to the vast and dynamic landscapes of stock markets. For example, Nayak et al. (2015) exploited machine learning algorithms based on manual indicators, but the underlying random walk hypothesis may hamper the task of understanding inherently non-stationary series. With the rise of different architectures, Feng et al. (2019) introduced Relational Stock Ranking (RSR) to catch the sentiment of the market in trading data.
@@
Although there exist several previous experiments with ML factors that attempted to fetch market sentiment in the quantitative finance field, this research still faces two difficulties. Firstly, in the field of sentiment analysis, which is a branch of Natural Language Processing (NLP, a field of computer science that aims to understand, interpret, and generate human language), models are used to convert words in text to word vectors through word embeddings to serve as inputs. However, instead of words, financial datasets contain both categorical data, such as industry types, as well as quantitative data, such as price fluctuation, turnover rate, and financial indicators. If the input comprises only categorical data, the time series can be treated as a sentence (Gorishniy et al., 2022). In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
Secondly, most NLP tasks can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition. As an example, the transformer architecture is based on the seq2seq architecture (Vaswani et al., 2017). To utilize existing outputs, decoders in transformer sequentially output samples and use masking operations to handle input sequences during training. However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
To address these problems, we propose quantformer, a modified transformer architecture adapted to quantitative data, and use it as an investment factor. Quantformer is able to input numerical data directly, which refers to a method similar to sentiment analysis. Our contributions lie in following aspects:
  • We propose quantformer, whose structure adapts to rolling stock-related time series data as inputs without positioning module and word embedding. The new structure with linear embedding shows better fitting to the numerical type inputs.
  • According to our experiments, quantformer-based factors perform better in back-tests compared with other traditional factor strategies under the same time period with different trading frequencies. These experiments indicate a potential direction for applying transformer-based models to quantitative financial tasks.
The paper is structured as follows. Section 2 discusses previous quantitative financial works based on machine learning. Sections 3 and 4 introduce quantformers. A factor based on quantformer will be trained and back-tested. For the practical backtest, we collected data from more than 4,600 stocks in the past 14 years (from 2010 to 2023). To comprehensively test the ability of the factor, we divide the data by different frequencies (Subsection 4.2) and trained under different training scales (Subsection 5.2). Finally, the results of the back-tests including the comparison between the quantformer-factor and other 100 factors as well as the insights gained from such comparative analysis are discussed in Sections 4 and 5.
@@ 2 Preliminaries
This section briefly introduces the related works about stock prediction with market sentiment and the development of quantitative financial trading methods, especially those based on Machine Learning (ML) methods.
@@ 2.1 Stock data and market sentiment
It is known that, in some cases, stock data can be deemed to influence the future trend of a stock price (Asness, 1995). Chen et al. (2020), P.H and Rishad (2020) have discussed the correlation between investor sentiment and stock market metrics such as turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns. This relationship is not uniform but varies with changes in investor sentiment. It also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions. This link between sentiment and market performance is further affirmed by studies focusing on factors such as the turnover ratio, which are seen as reflective of market liquidity and investor behavior (Naseem et al., 2021). Ding et al. (2023) built a new factor based on sentiment analysis with the average of trading signals from technical trading strategies to benchmark stocks of the S&P 500 index and DJIA. The sentiment factor shows correlation with stock returns. These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
@@ 2.2 Applications of ML methods for trading
2.2.1 Support Vector Machines (SVMs)
SVMs have been used in quantitative trading, as they can manage high-dimensional spaces and intricacies found in datasets. Kim (2003) used SVMs to forecast financial time series in 2003, with a focus on the South Korean stock market, exhibiting the robustness of SVM against traditional statistical methods and some other computational techniques. Similarly, Huang et al. (2005) showcased SVMs’ prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments. Cavalcante et al. (2016) assessed SVMs’ adaptability and precision in financial tasks such as portfolio and quantitative trading.
However, SVMs can also show some disadvantages in quantitative trading. In Huck (2009)’s work, SVM was underscored for the accuracy of overfitting. When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, making the strategy ineffective on unseen data. In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptations (Kercheval and Zhang, 2015).
@@ 2.2.2 Long Short-Term Memory (LSTM)
LSTM is a neural network architecture designed to capture long-term dependencies in sequential data (Hochreiter and Schmidhuber, 1997). LSTM networks can accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models. Orsel and Yamada (2022) combined LSTM with linear Kalman filter and set experiments on different volatility stocks to build strategies. Bao et al. (2017) deployed LSTM with stacked autoencoders to forecast financial time series. Fischer and Krauss (2018a) tested the efficacy of LSTM for price predictions. They observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over other conventional models. Sezer and Ozbayoglu (2018) combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy. For predicting stock prices, Zhang et al. (2017) harnessed the power of LSTM to unearth multi-frequency trading patterns. Their method underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
@@ 2.2.3 Gated Recurrent Units (GRUs)
GRUs (Chung et al., 2014) are another neural network architecture which has emerged as a noteworthy method in the quest for accurate financial forecasting. GRUs can capture long-term dependencies in time series data and have seen increasing popularity in the world of quantitative trading. Bao et al. (2017) introduced a comprehensive deep-learning framework for financial time series, which incorporated GRUs alongside LSTM and stacked autoencoders. Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs. Fischer and Krauss (2018b) centralized on LSTM networks, provided comparative insights into GRUs. However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints. Beyond that, Pak et al. (2018) presented a model fusing GRUs with Convolutional Neural Networks (CNN) O’Shea and Nash (2015). The strategy to combine CNNs and GRUs also exhibits advantages for stock price prediction. Sun et al. (2024) introduced the Attention-GRU model which combined attention and GRU to establish a new factor based on CVaR portfolio (Zhu and Fukushima, 2009, Ban et al., 2018). The Attention-GRU model fitted market return, stocks return from 28 Dow Jones Industrial Average index (DJIA) stocks, which achieved better performance than other models in eight metrics such as annual return, standard deviation and information ratio.
2.3 Transformer in time series prediction
Transformer (Vaswani et al., 2017), which was introduced in 2017, has received considerable attention in the NLP field. According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models. Transformer-based models such as GPT (Radford et al., 2019) and BERT (Devlin et al., 2019) have been influential in the NLP field. More recent transformer-based models include ChatGPT (Brown et al., 2020), FinBERT (Araci, 2019, Yang et al., 2020) and BloombergGPT (Wu et al., 2023).
Similar to other neural sequence transduction models, transformer also has an encoder-decoder structure. The model takes as input a sequence of context, which could be natural language sentences, and generates output with texts depending on the input, similar to a translation according to the input or a predicted sentence. Transformer consists of stacks of layers, which include self-attention and point-wise operations. These layers are used to connect the encoder and decoder in the architecture. For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation. The linear and softmax functions are used in the dot-product attention module. In the vanilla transformer model, multi-head attention layers are used and the result from each layer is outputted after going through an additional normalization layer. The above-mentioned steps cover most of the progress made with the vanilla transformer.
Table 1 shows a set of advanced transformer models, with a focus on time series prediction tasks, which are tasks of interest in quantitative finance. The table highlights the modifications made to the original transformer model. These modifications can be categorized into innovations in attention mechanisms, encoding and decoding structure, positioning module, and mixture model.
Some research proposed enhancements of the attention module. For example, Adversarial Sparse Transformer (AST) (Wu et al., 2020) introduced sparse attention using the $\alpha$-Entmax function within adversarial training. Autoformer (Wu et al., 2021b) replaced traditional self-attention with an auto-correlation mechanism that utilizes periodic patterns in frequency-domain time series analysis. Similarly, Informer (Zhou et al., 2021) suggested ProbSparse attention that selects only the most relevant queries and introduces generative-style decoding to handle long-term dependencies. Gated Transformer Networks (GTN) (Liu et al., 2021) used separate attention for channel-wise and temporal-wise interactions for multivariate time series. Patch Time Series Transformer (PatchTST) (Nie et al., 2022) divided time series into patches to model local dependencies, and Quatformer (Chen et al., 2022) attempted to use rotation-based attention to capture periodic patterns in time series. AirFormer (Liang et al., 2023) introduced dartboard spatial attention and causal temporal attention in air quality prediction. Rough transformer (Moreno-Pino et al., 2024) facilitated multi-view signature attention and continuous-time path encoding for financial market modeling for complex financial time series prediction.
In another direction, some works tried to make changes to the encoding or decoding of transformer. FEDformer (Zhou et al., 2022) splitted time series data into trend and seasonal components with frequency-enhanced modules. To improve computational efficiency, some models, including Multivariate Transformer (Zerveas et al., 2021) gave up using a decoder and focused solely on an encoder for feature extraction for forecasting tasks.
As a kind of revision sight, some research has proposed to replace or extend the positioning modules to better accommodate time series data. For instance, HFformer (Barez et al., 2023) gave up positioning modules, and uses activation functions with a lightweight linear decoder to revise the model. Noting that temporal order is inherently embedded in time series data, iTransformer (Liu et al., 2024) introduced a dimension-inverted structure and removes positioning. Transformer Hawkes Process (THP) (Zuo et al., 2020) put time intervals into the positioning module directly. FX-Spot Transformer (Fischer et al., 2024) used the Time to Vector (Time2Vec) function to handle irregular time intervals.
To overcome the limitations of standalone transformer models, an approach is to integrate them with other frameworks. ConvLSTM (Kim et al., 2024) merged convolutional neural networks (CNNs) and LSTM layers with transformer moduled to simultaneously model spatial and temporal patterns. ProTran (Tang and Matteson, 2021) integrated state-space models (SSMs) with transformers to address non-Markovian dynamics in time series data, providing robust long-term forecasts. These hybrid approaches improved the versatility and domain adaptability of transformer-based architectures.
Table 1: Advanced transformer-based architectures for time series data
Reference Model Method
Wu et al. (2020) Adversarial Sparse Transformer (AST) Sparse attention with adversarial training
Liang et al. (2023) AirFormer Dartboard spatial self-attention and causal temporal self-attention for different training stage
Wang et al. (2022b) ALSP-TF Local interaction and global interaction between layers
Wu et al. (2021b) Autoformer Auto-correlation instead of attention
Wang (2023) BiLSTM-MTRAN-TCN Combined BiLSTM, Transformer and Temporary Revolution Network (TCN)
Kim et al. (2024) ConvLSTM Combined LSTM with transformer
Zeng et al. (2023a) LTSF-Linear DLinear and NLinear, MLP-based with transformer
So et al. (2019) Evolved Transformer Neural Architecture Search (NAS) in transformer
Zhou et al. (2022) FEDformer Timestamp encoding with frequency enhanced attention
Liu et al. (2021) Gated Transformer Networks (GTN) Two-tower transformer with channel-wise & step-wise
Ding et al. (2020) Gaussian transformer Multi-scale Gaussian prior, orthogonal regularization and trading gap splitter for self-attention
Barez et al. (2023) HFformer Linear decoder and spiking activation
Zhou et al. (2021) Informer ProbSparse self-attention, self-attention distilling operation and generative style decoder
Liu et al. (2024) iTransformer Dimension inversion and deleted positioning
Li et al. (2019) LogSparse Transformer Causal convolution attention and query-key matching
Eisenach et al. (2020) MQ-Transformer Positional encoding from event indicators and decoder-encoder attention for context-alignment
Ramos-Pérez et al. (2021) Multi-Transformer Average multi-head and bagging attention and positioning based on time stamps
Zerveas et al. (2021) Multivariate transformer Deleted decoder and applied learnable positioning module for multivariate time series
Liu et al. (2022b) Non-stationary Transformer Series standardization and De-stationary attention
Nie et al. (2022) Patch Time Series Transformer (PatchTST) Divided time series into patches, applied channel-independence and instance normalization
Tang and Matteson (2021) ProTran Modeled non-Markovian dynamics by State Space Models (SSMs) to
Liu et al. (2022a) Pyraformer Pyramidal Attention Module (PAM) and intra-scale neighboring connections model
Chen et al. (2022) Quatformer Learning-to-rotate Attention (LRA) with global memory and trend normalization
Kitaev et al. (2020) Reformer Locality-sensitive hashing attention and reversible residual layers
Moreno-Pino et al. (2024) Rough Transformer Multi-view signature attention and continuous-time paths encoding transformation
Shabani et al. (2022) Scaleformer Multi-scale framework with various models (such as FEDformer, Autoformer)
Lin et al. (2021) State Space Decomposition Neural Network (SSDNet) Transformer with SSMs
Ma et al. (2025) Stockformer Dual-frequency spatiotemporal encoder for data in different frequency with graph embedding and fusion attention
Chowdhury et al. (2022) TARNet Attention score for timestamps masking
Olorunnimbe and Viktor (2024) Temporal Fusion Transformer (TFT) LSTM learning positional encoding
Fischer et al. (2024) Transformer for FX-Spot Time2Vec positioning module
Zuo et al. (2020) Transformer Hawkes process (THP) Positional encoding by translating time intervals
Cirstea et al. (2022) Triformer Triangular,variable-specific patch attention